OpenStudy (anonymous):

what is the formula "in four hours a hiker goes from 892 feet to 256. what is the rate of speed"

hero (hero):

\[r_s=\frac{d}{t} = \frac{892 - 256}{4} \dot\ \frac{\text{ft}}{\text{hr}}\]

OpenStudy (anonymous):

thanks

hero (hero):

You're welcome